OpenStudy (anonymous):

find an ordered triple to represent x in x=3y-z if y=<2,-1,5> and z=<1,-6,6>

OpenStudy (perl):

did you plug it in

OpenStudy (anonymous):

do i add it or what?

OpenStudy (perl):

$$ \Large x = 3~\cdot <2,-1,5> - <1,-6,6>\\ \Large = <3 \cdot 2, 3 \cdot(-1), 3 \cdot 5> - <1,-6,6>\\ \Large = < 6, -3, 15> - <1,-6,6>\\ $$

OpenStudy (perl):

@SolomonZelman I find that negative sign to be too big. do you know latex to shrink that

OpenStudy (anonymous):

so it would be 5,3,9?